A bottle dropped from a balloon reaches the ground in 20.0 s. Determine the height of the balloon if a) it was at rest in the air and b) it was ascending with a speed of 50.0 m/s when the bottle was dropped


1
Expert's answer
2022-03-04T08:31:50-0500

a)

"h=\frac{gt^2}2=2~km,"

b)

"h'=\frac{gt^2}2-vt=1~km."